Please answer all the questions! Thanks!

Toys for fun Co. is working at full production capacity producing 20,000 units of a unique product, OB1. Sale price and costs per unit for OB1 are as follows $ 40 Sale price Direct materials Direct manufacturing labour Manufacturing overhead Selling costs (all variable) $8 14 The unit manufacturing overhead cost is based on a variable cost per unit of $10 and fixed costs of 80,000 (at full capacity of 20,000 units). A Chinese customer, Zhao Co., has asked Toys for fun Co. to produce 4,000 units of Jedi1, a modification of OB1. Jedi1 would require the same manufacturing processes as OB1. Zhao Co. has offered to pay Toys for fun Co. $33 for a unit of Jedi1 and half the selling costs per unit. REQUIRED: Compute the contribution margin per unit for OB1 and Jedi1 What is the opportunity cost to Toys for fun Co. of producing the 4,000 units of Jedi1? (Assume that no overtime is worked) Should Toys for fun Co. accept the special order from Zhao Co.? Justify your recommandation Ottawa Co. has offered to produce 4,000 units of OB1 for Toys for fun Co. so that Toys for fun Co. may accept the Zhao Co. offer. That is, if Toys for fun Co. accepts the Ottawa Co. offer, Toys for fun Co. would manufacture 16,000 units of OB1 and 4,000 units of Jedi1 and purchase 4,000 units of OB1 from Ottawa Co. Ottawa Co. would charge Toys for fun Co. $30 per unit to manufacture OB1. Should Toys for fun Co. accept the Ottawa Co. offer? (Support your analysis with detailed analysis) Suppose Toys for fun Co. had been working at less than full capacity, producing 16,000 units of OB1 at the time the Zhao Co. offer was made. What is the minimum price Toys for fun Co. should accept for Jedi1 under these conditions (ignore the previous $33 selling price). Are there any qualitative factors that Toys for fun Co. should consider before making any decision on the special order from Zhao Co.? Discuss at least three factors. 1. 2. 3. 4. 5. 6.
Answer #1

Answer :

Calculation of Contribution Margin per unit of OB1 and Jedi1 (Calculation in $)

Particulars OB1 JEDI1

Selling Price per unit 40 37 (33+4)

Variable Cost per unit

Direct Material 5 5

Direct Labour 8 8

Variable Manu. Overhead 10 10

Variable Selling Overhead 8 8

Total Variable cost 31 31

Contribution Margin PU 9 6

Part 2 : Oppurtunity Cost

Oppurtunity Cost is the contribution foregone due to accepting Zhao's offer :

= Contribution PU * Units

= 9 * 4000 = 36000

Part 3 :

As per calculation shows above in Part1, contribution per unit (at full production capacity of OB1) is more than the contribution per unit (if XZhao's offers accepted).

So, there will be a loss of $3 contribution per unit if Zhao's offer accepted i.e. total loss of $12000.

So it is recommended to not to accept Zhao's offer and working at full production capacity of 20000 units of OB1.

Part 4.

Calculation if Ottawa's offer accepted by Toys for Fun co. (Calculation in $)

Particulars OB1 (Manu) OB1 (Pur) JEDI1 Total

Contribution Margin PU 9 10 (40-30) 6

No. Of Units 16000 4000 4000

Total Contribution 144000 40000 24000 208000

Fixed Cost 80000

Profit 128000

Calculation of Profit in Normal case

Contribution = 20000 * 9 = 180000

Fixed Cost = = 80000

Profit = 100000

So Toys for fun co. should have to accept Ottawa's offer as it will increase their profit by $28000.